OpenStudy (anonymous):

simplify complex fractions:

OpenStudy (amistre64):

complex as in with the imaginary "i"?

OpenStudy (anonymous):

all you need to know is i^2 = -1

OpenStudy (anonymous):

no mutiplying and dividing rational expressions \[(x^3y^2z)/(a^2b^2)/(a^3x^2y)/(b^2)\]

OpenStudy (anonymous):

where the slash is its written like a fraction

OpenStudy (anonymous):

So what is over what here?

OpenStudy (anonymous):

stack the parenthisis up ontop of each other where it looks like a fraction with things. idk how to punch it in on here were it looks like it does on my worksheet

OpenStudy (amistre64):

you can use // to mean the main divider bar between them, or some other version like */* just explain the symbol :)

OpenStudy (anonymous):

From what it looks like, I would assume it's like a rational expression over another. Correct?

OpenStudy (amistre64):

the b^2.s cancel out

OpenStudy (anonymous):

yes rational expressions

OpenStudy (amistre64):

basic strategy as with all fraction on fraction math is to turn the bottom one upside down and multiply across... 1/5 // 6/3 = 1/5 * 3/6 = 3/30

OpenStudy (anonymous):

but when its all letters and exponets how do you do it

OpenStudy (amistre64):

you mean the multiply part? or the flipping part?

OpenStudy (anonymous):

the multiply part. flipping i get now but idk what to do with the letters

OpenStudy (amistre64):

make it easier on yourself and split them up... a^4 means aaaa c^2 mean cc r^7 means rrrrrrr

OpenStudy (amistre64):

now, x^3 times x^4 means xxx xxxx = xxxxxxx = x^7

OpenStudy (amistre64):

c^4 times c^5 means cccc ccccc = ccccccccc = c^9

OpenStudy (anonymous):

yea i get the x^4 means xxxx but what about x^3y^2z

OpenStudy (amistre64):

xxx yy z is all that is

OpenStudy (amistre64):

squish em all together, then count them back out again....

OpenStudy (anonymous):

what your confusing me

OpenStudy (anonymous):

a/x = a* 1/x, therefore \[(x ^{3}y ^{2}z)/(a ^{2}b ^{2})/(a ^{2} x ^{2} y)/(b ^{2})] = \[(x ^{3} y ^{2} z)/(a ^{2} b2) * 1/(a ^{2} x ^{2} y)/(b ^{2})] = \[(x ^{3} y ^{2} z)/(a ^{2} b ^{2})*(b^{2})/(a ^{2} x ^{2} y)\]

OpenStudy (amistre64):

xxx yy z bb -------- x ------ aa bb aaa xx y

OpenStudy (amistre64):

anything that is the same on the top and bottom can be tossed out, it just means it equals 1. x y z 1 x y z xyz ----- x ---- = ------ = ---- aa aaa aaaaa a^5

OpenStudy (anonymous):

still not makin a whole lota sense

OpenStudy (amistre64):

which part is not making sense?

OpenStudy (amistre64):

i flipped it, you said thatpart you understood.... whats left?

OpenStudy (amistre64):

Does: xx y bb ------ = 1? xx y bb

OpenStudy (anonymous):

the whole thing. its a fraction placed ontop on a fraction saying simplify. like am i supposed to add the exponets together, flip them and multiply, or flip and divide

OpenStudy (amistre64):

flip the bottom, thats always the first step. split the exponents up so you can see what you got whatever is the same from top to bottom cancels out to 1 squish the rest together and number them again with an exponent.

OpenStudy (amistre64):

lets keep it simpler, how would I start this? x^2/y^3 // y/x^2

OpenStudy (anonymous):

flip it and take away the x^2 and just leave it as xy^3/xy i think

OpenStudy (amistre64):

one step at a time, dont rush ahead... lets flip the bottom x^2 x^2 ---- * ---- now what do I do? y^3 y

OpenStudy (anonymous):

the x^2 ---- y part

OpenStudy (anonymous):

oh wait y^3 x y ------- x^2 x^2

OpenStudy (amistre64):

you let me worry about making it look good on the screen, just tell me what we do next...

OpenStudy (amistre64):

I already flipped it, x^2 x^2 ---- * ---- now what do I do? y^3 y

OpenStudy (anonymous):

i want to say crossmultiply

OpenStudy (anonymous):

but thats usually not with letter

OpenStudy (amistre64):

we do that if there is an (=) between them, there is no (=) here so we dont "cross" multilply, we just multiply straight across

OpenStudy (anonymous):

ok just add the exponets together

OpenStudy (amistre64):

-------------->> x^2 x^2 x^2 x^2 ---- * ---- = ---------- y^3 y y^3 y -------------->> does this look right to you?

OpenStudy (anonymous):

not really shouldnt it be x^3/y^4

OpenStudy (amistre64):

not yet.... but does the setup look right? did I multiply it across correctly?

OpenStudy (anonymous):

yea i think.

OpenStudy (amistre64):

now lets add exponents to get the final result x^2 x^2 x^(2+2) x^4 ------- = --------- = ---- y^2 y y^(2+1) y^4

OpenStudy (amistre64):

2+1 = 3... sorry, forgot how to add :)

OpenStudy (anonymous):

that looks more like it

OpenStudy (amistre64):

so lets try your original problem and see if we can step thru it ok?

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

x^3 y^2 z --------- a^2 b^2 --------------- a^3 x^2 y -------- b^2 our first step is to do what?

OpenStudy (anonymous):

im working it on a sheet of paper and the first thing i did was put the bottom fraction beside it and put a division sign between them

OpenStudy (amistre64):

ok.... x^3 y^2 z a^3 x^2 y --------- / ---------- a^2 b^2 b^2 now what?

OpenStudy (anonymous):

um flip the entire thing upside down or add the exponets

OpenStudy (amistre64):

only flip the right side fraction... not the "whole" problem....JUST that right side gets flipped. right?

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

just like this.... x^3 y^2 z b^2 --------- ---------- a^2 b^2 a^3 x^2 y

OpenStudy (anonymous):

yea then add the exponets together

OpenStudy (amistre64):

what do you mean by add the exponents together? that is not really something that needs to be done just yet. Look at the top and the bottom of this oversized fraction and see if we can cross out stuff that looks the same.

OpenStudy (amistre64):

lets "line" things up from top to bottom....are we allowed to do that?

OpenStudy (anonymous):

line things up idk what your talking about

OpenStudy (amistre64):

I am going to move the stuff around so that it looks better. Like this..... x^3 y^2 z b^2 -------------------- x^2 y a^2 a^3 b^2

OpenStudy (amistre64):

what can you see that we can "get rid of" that is the same from the top and the bottom?

OpenStudy (anonymous):

um on the bottom combine the a^2 and a^3 together to get a^5

OpenStudy (amistre64):

good we can do that... x^3 y^2 z b^2 -------------------- x^2 y a^5 b^2

OpenStudy (anonymous):

can we also combine the x^3 and x^2 or not since there on seperate parts of the equation

OpenStudy (amistre64):

we cant "add" them together but watch this: xx x ---- = what? xx

OpenStudy (anonymous):

5

OpenStudy (amistre64):

no....not quite. do you remember that anything when placed over itself is equal to 1? 2 -- = 1 2 6m --- = 1 6m xx -- = 1 xx

OpenStudy (anonymous):

i think

OpenStudy (amistre64):

this is a fundamental concept in math. whenever we "divide" a number by itself we get an answer of 1 1 --- 8 | 8 right?

OpenStudy (anonymous):

i guess but i dont particularly see how that would work in this type of problem

OpenStudy (amistre64):

all fractions are is division. what is 1 divided by 2 = 1/2 3 divided by 5 = 3/5 15 divided by 5 = 15/5 all fractions are division......

OpenStudy (amistre64):

when we divide a number by itself we get 1 9 divided by 9 = 9/9 = 1

OpenStudy (anonymous):

but in this case were dividing x^3/ x^2

OpenStudy (amistre64):

exactly :) which is why I like to split it up so that you can "see" what is going on with exponents. x x x xx x ---- = --- -- = 1x = x x x xx 1

OpenStudy (anonymous):

so like that i can just "get rid of" the x part in the equation

OpenStudy (amistre64):

not the whole thing... but you can get rid of alot of x s how many x's do I have left in my solution up there?

OpenStudy (anonymous):

the x^3 and the x^2 so 5 i think

OpenStudy (amistre64):

x x x xx x ---- = --- -- = 1x = x x x xx 1 ^ how many x's do I have right here?

OpenStudy (anonymous):

1

OpenStudy (amistre64):

then we are left with 1 lonely little x that we have to leave in this equation.... right?

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

you know already that we we multiply exponents we "add" then together. When we do the opposite of multiplication (which is division) we do the opposite to the exponents too. we "SUBTRACT them. x^3/x^2 = x^(3-2) = x^1 = x

OpenStudy (anonymous):

ok so how do we plug this one lonely x into the equation

OpenStudy (amistre64):

we just put it back were we got it from :) like this: x y^2 z b^2 ------------- y a^5 b^2

OpenStudy (anonymous):

ok what next. same thing to the b^2s

OpenStudy (amistre64):

yep, that would be good. b^2/b^2 is division.... so we SUBTRACT exponents. b^(2-2) = b^0 = 1 which just disappears

OpenStudy (anonymous):

ok so the b is just slapped onto the end of the top part of the equation without the exponent on it

OpenStudy (amistre64):

there is no b to put back if b^3 means bbb and b^1 means b b^0 means _______ no little b's to do anything with, zero b's, they are gone......

OpenStudy (anonymous):

so on the actual equation after this last part the b will be no more

OpenStudy (amistre64):

thats correct, they vanished...... x y^2 z --------- now whats left to work with? y a^5

OpenStudy (anonymous):

the y

OpenStudy (amistre64):

very good :) so, can you show me how we work that out?

OpenStudy (anonymous):

y^2/y^19(which is just nothing) subtract them and it leaves y

OpenStudy (amistre64):

perfect!!

OpenStudy (amistre64):

x y z ------ we appear to have come to the end :) a^5

OpenStudy (anonymous):

that 9 wasnt supposed to be there my finger sliped off the shift button when i hit the paraenthasis